# Division of labor enables efficient collective decision-making under uncertainty

### v1

@@ Abstract
How do social animals make effective decisions in the absence of a leader? While coordination can improve accuracy, it also delays responses as information propagates through the group. In changing environments, these delays can outweigh the benefits of centralized control, making decentralized strategies advantageous in large groups. This raises a key question: how can groups implement efficient collective decisions without central coordination? We address this question using a model of collective foraging in which individuals choose whether to invest in costly exploration or remain idle, while sharing information and rewards across the group. We show that decentralized collectives can match the performance of centrally controlled groups through a division of labor: a small, heterogeneous subset explores even when expected rewards are negative, while a synchronized majority forages only when expected rewards are positive. Information redundancy causes the optimal scout number to grow sublinearly with group size, so that larger groups need proportionally fewer explorers. The heterogeneity of the group is maximized at intermediate ecological pressures, but optimal groups are homogeneous when costs or environmental contrasts or fluctuations are extreme. Crucially, these group-level policies do not require central coordination, emerging instead from agents following simple threshold-based decision rules. We thus demonstrate a mechanism through which leaderless collectives can make effective decisions under uncertainty and show how ecological pressures can drive changes in the distribution of strategies employed by the group.
@@ 1 Introduction
One solution to this problem is centralized leadership. In many wild wolf packs the breeding pair tends to initiate movements and coordinates traveling, resting, and hunting [Mech1999]. Human history offers parallel examples. Admiral Zheng He commanded fleets across the Indian Ocean in the early fifteenth century on imperial orders, spurring Chinese emigration and trade throughout Southeast Asia [Levathes2014]. A century later, royal sponsorship supported Magellan’s circumnavigation, which opened the first western sea route to the Spice Islands and accelerated global trade and contact between distant regions [Bergreen2003]. In each case, a designated explorer, acting under a central authority, bore enormous personal risk, while the nations that sent them reaped lasting rewards. However, most social animals including honeybees, ants, and fish schools lack fixed leaders and coordinate through local interactions [Pratt2002, Seeley2004, Couzin2005, Sumpter2005, Seeley2011, Hein2015]. Despite extensive empirical study of decentralized systems [Pratt2002, Sumpter2005, Giardina2008, Gordon2014], it remains unclear how closely decentralized coordination matches centralized control, how performance depends on diversity in individual strategies, and under what conditions heterogeneous strategies offer an advantage in leaderless groups.
@@
To address these questions we develop and analyze a stochastic foraging model [Puterman1994, Mahadevan1996, Gold2007, Kilpatrick2021, BlumMoyse2025, Dembo2010, Bressloff2021, Barendregt2025] to understand how collectives can achieve efficient exploration. We show that well-calibrated collectives exhibit characteristic dynamics: during times of plenty, risk-averse individuals venture out together, amplifying collective gains. When resources are scarce, most individuals remain idle. Yet, a small subset of risk-tolerant explorers continues foraging even when expected rewards are negative, and can rapidly detect environmental improvements. Composition is critical as population dominated by cautious individuals fail to discover new resources, while those dominated by risk-tolerant foragers exhaust shared reserves. Indeed, the number of risk-tolerant individuals grows sublinearly with population size, making a balanced division of labor between exploration and exploitation essential for efficient foraging [Seeley2011, Charbonneau2015]. Remarkably, collectives with a well-calibrated mixture of cautious and bold individuals can perform as well as groups governed by a central decision-maker. The most heterogeneous compositions emerge under intermediate ecological pressures, when exploration costs and environmental variability are balanced. We thus show how simple individual decision rules lead to optimal collective performance without coordination, and predict when ecological pressures favor heterogeneous versus uniform group composition.
@@
Figure 1: Overview of model and setup. Individuals in a collective observe returning foragers to update their belief about the environmental state and select actions. (A) The belief, $g^{t}$, is updated based on the observed outcomes of foraging attempts. (B) Individuals forage when their belief that the environment is in a high-reward state exceeds a threshold, $\theta$; these thresholds can differ between individuals. (C) Foraging is successful with probability $\gamma^{t}$, which switches between high and low values with the environmental state. (D) Under centralized control, a planner decides on the optimal number of foragers to maximize returns. (E) Under decentralized control, coordination emerges from individual foraging decisions.
@@ 2 A Model of Collective Foraging under Uncertainty
Collective decisions about foraging in large, decentralized collectives face a fundamental challenge: How do groups balance individual risk and collective benefit in dynamically changing environments without central coordination? To address this question we consider a minimal model that shows that decentralized decision-making mechanisms can yield optimal collective long-term returns, even when individual expected returns are negative. Here we provide an overview of the model; mathematical details can be found in SI Appendix, section 1.
@@
We assume that the community coordinates its behavior through shared information about the outcomes of foraging attempts. Returning foragers are observed by the entire community, so all individuals know how many foraging attempts have been successful (Fig. 1A). Individuals are Bayesian observers [Wald1948, Bogacz2006, VelizCuba2016] and compute the probability that the environmental state is in a high-reward state based on the observed number of successful foragers and knowledge about the environment’s volatility. We refer to this subjective probability as their belief about the environmental state. Because all individuals have access to the same information, they share a common belief that guides their foraging decisions.
@@
Maximizing returns requires optimizing forager number: too many foragers when conditions are poor wastes effort; too few when conditions are good misses opportunities. To establish an upper bound on collective performance, we derive a theoretical benchmark using a centralized planner that optimally determines foraging numbers to maximize long-term collective returns (Fig. 1D). Under this centralized policy, all individuals follow the directives of a planner. This planner infers the environmental state by observing foraging successes to determine the best number of foragers. This optimal policy can be computed using dynamic programming [Bellman1954, Mahadevan1996, Sutton1998] and provides an upper bound on achievable performance. We compare this benchmark with outcomes from independent agents who make individual decisions based on shared information but without central coordination.
@@ 3 Results
We use our model of foraging and belief-updating to ask how a collective translates shared information about the environment into an effective foraging decisions. We begin by establishing a normative benchmark defined by the allocation selection of a centralized decision-maker to maximize expected payoff. We then show that a leaderless group can achieve optimal performance using heterogeneous individual threshold rules, and analyze how environmental conditions shape structured heterogeneity in an optimal, decentralized collective.
@@ 3.1 Robust collective coordination through structured heterogeneity
A central decision-maker can select an optimal number of foragers to maximize expected payoffs on each turn. In a decentralized collective, individuals share a common belief about the environmental state but make individual choices based on a private threshold. If all individuals choose a threshold that maximizes their own expected return, then all thresholds are identical. Since they share the same belief, everyone makes the same decision: forage when the expected reward is positive, and remain idle when the expected reward is negative. As a result, a collective can quickly detect when an environment switches from a high- to a low-rewarding state. However, when the collective believes the reward state is likely to be low, no individuals forage, and the collective will not detect a switch back to a high-rewarding state. The resulting failure to explore thus leads to starvation even if good times have returned. Conversely, if all individuals adopt low thresholds and forage often, the collective can quickly infer the environmental state but incurs large energetic losses in lean times. Effective collective performance therefore requires a division of labor between a minority that forages even when rewards are unlikely, and a majority that waits for favorable conditions.
@@
With an appropriate distribution of individual decision thresholds, a decentralized collective can achieve the same expected returns as an optimally controlled group. We show this by constructing a threshold distribution that matches the number of foragers deployed under the optimal centralized policy. The optimal strategy of a central planner can be represented by a function, $\pi(g),$ that maps the collective belief about the state of the environment, $g,$ to the number of individuals sent out to forage, $n=\pi(g)$ (Fig. 2A). As the belief that the environment is in a high-reward state increases, the planner commits more foragers [Puterman1994, Topkis1998]. The policy defines a sequence of belief thresholds, $\theta_{n}=\inf\{\theta\in[0,1]:\pi(\theta)\geq n\}$ for $n=1,2,\cdots,N$, where $N$ is the group size, so that multiple thresholds can be the same. Each threshold, $\theta_{n},$ is the smallest belief value at which sending out exactly $n$ foragers is optimal. In particular, the majority of thresholds are the same at the critical belief $\theta_{c}$ (Fig. 3), at which the expected foraging return is zero.
@@
Figure 2: Division of labor enables centralized and decentralized coordination. (A) In decentralized collectives, individuals have different foraging thresholds. Risk-averse “deliberators” share a common threshold, $\theta_{c},$ and forage only when the expected reward is positive. Risk-tolerant “scouts” have lower, heterogeneous thresholds and forage even when expected rewards are negative. (B) Under centralized control, a planner implementing the optimal policy, $\pi(g),$ deploys an increasing number of foragers as the belief increases. An appropriate distribution of individual thresholds allows a decentralized collective to match this allocation (compare panels A and B). (C) The optimal number of risk-tolerant scouts, $N_{s},$ grows sublinearly with population size, NN, so an increasingly small fraction of explorers suffices for efficient foraging. (D) The collective belief, $g^{t},$ tracks environmental fluctuations; scouts continue exploring during lean periods, enabling rapid re-engagement when conditions improve.
@@
The strategy of a decentralized collective is fully determined by a set of private thresholds. Hence, a collective of size $N$ whose individuals have thresholds $\{\theta_{n}\}_{n=1}^{N}$ will send out exactly $n$ foragers when the belief satisfies $\theta_{n}\leq g<\theta_{n+1}$, with the convention $\theta_{N+1}=1$ (Fig. 2B). This choice of thresholds thus results in a number of foragers that equals that dictated by the optimal policy, $\pi(g)$ for any belief $g$. Importantly, this is a consequence of individual decisions based on private thresholds, rather than orders from a central planner. Such a distribution of foraging strategies could arise from evolved or learned differences in risk sensitivity [Seeley1983, Dreller1998, Liang2012], consistent with classic response-threshold models of division of labor [Bonabeau1996, Beshers2001].
@@
Collectives governed by a central decision-maker are inherently vulnerable to the loss or failure of this planner. By contrast, decentralized collectives with heterogeneous individual strategies and are therefore more likely to be evolutionarilare more can be more robust to perturbations, which may promote evolutionary stability [Jandt2013, Couzin2005]. As long as phenotypic differences across the population generate an appropriate distribution of risk tolerances, the collective can achieve optimal foraging performance [Seeley1983, Charbonneau2015]. Structured heterogeneity allows decentralized groups to match the allocations of a central planner, yielding robust coordination and sustained performance in dynamic environments. We next analyze the structure of the optimal population, and ask how ecological pressures shape this distribution of strategies.
3.2 Exploratory scouts and synchronized deliberators
@@
Figure 3: Effects of perturbing the distribution of decision thresholds. (A) Optimal threshold distribution for a collective of $N=100$ individuals. (B-D) Perturbed strategy distributions: (B) desynchronizing the risk-averse majority by spreading their thresholds; (C) synchronizing risk-tolerant individuals at a high threshold; (D) synchronizing scouts at a low threshold. Colors indicate strategy types. (E) A trade-off plot of return and accuracy for the optimal strategy and perturbations shows that returns are maximized at intermediate accuracy (gray dot; panel A), reflecting weakened coordinated exploitation under desynchronization (brown–gray curve; panel B), improved accuracy at increased cost under moderate heterogeneity (pink–gray curve; panel C), and degraded returns under excessive risk tolerance (blue–gray curve; panel D). Returns are normalized by the maximum possible return $\rho_{\max}=(\gamma_{+}-\lambda)/2$, achieved by committing the entire collective only in the high reward state.
@@
We next determine the optimal balance between risk-tolerant individuals who forage under uncertainty and risk-averse individuals who forage only when the expected rewards is positive. Empirical studies show a consistent pattern in many social species: a small fraction of individuals, such as “scouts" in honeybee populations, persistently engages in exploration, while most colony members adopt safer strategies and exploit known resources [Seeley1983, Seeley2011, Beekman2001, Dornhaus2006]. We show that the optimal decentralized strategy yields a similar pattern. To do so we use the critical threshold $\theta_{c}$ and classify individuals with decision thresholds below $\theta_{c}$ are risk-tolerant, and those with thresholds equal to or exceeding $\theta_{c}$ as risk-averse.
@@
In small optimal collectives, all individuals are risk-tolerant, because each foraging outcome is highly informative and has a large impact on the collective belief about the environmental state (Fig. 2C). As group size increases, performance improves, but the optimal division of labor changes. The expected cost of foraging grows linearly with the number of active foragers, whereas additional foraging outcomes provide redundant information and thus become less valuable. A mathematical analysis shows that the optimal number of risk-tolerant individuals grows logarithmically with group size (dashed line in Fig. 2C). This sublinear scaling is robust across a wide range of environmental volatilities, foraging costs, and reward structures (SI Appendix, Fig. S1). Consequently, larger groups comprise an increasing number of risk-averse individuals, and a correspondingly smaller fraction of risk-tolerant explorers that achieve efficient collective performance. This prediction is consistent with empirical and theoretical studies showing that effective animal collectives rely on a small exploratory minority whose size grows more slowly than group size [Couzin2005, Cronin2014]. Details of the asymptotic analysis underlying this result are provided in SI Appendix, section 2.
@@
While the decision thresholds of risk-tolerant individuals are distributed, all risk-averse individuals share a common threshold, $\theta_{c}$. Indeed, the entire collective commits to foraging once the expected reward becomes positive (see SI Appendix and Fig. S2). This synchronization is a consequence of the linear increase of total reward with the number of committed individuals, implying that full commitment is optimal when returns are expected to be favorable. Thus, an optimal collective consists of a heterogeneous minority that maintains environmental awareness though continued sampling, and a synchronized majority that maximizes collective returns through coordinated exploitation once conditions are deemed favorable.
@@
Optimal collective foraging strategies exhibit a characteristic cascade of activity in response to environmental changes (Fig. 2D). When conditions deteriorate, widespread foraging produces few successes, quickly shifting the collective belief toward low expected rewards. Most individuals stop foraging, leaving a small subset of risk-tolerant explorers that continue sampling. When conditions deteriorate severely, even limited exploration ceases to minimize costs. Nevertheless, in the absence of new information, collective belief gradually relaxes toward uncertainty, prompting renewed exploration. If some of these sparse foraging attempts are successful, the belief shifts towards high expected rewards, and more foragers are recruited. Once the collective belief crosses the threshold $\theta_{c}$, coordinated collective foraging is triggered again, and the process repeats. Heterogeneity in decision thresholds allows the population to act as a distributed sensor, enabling accurate tracking of environmental changes while optimizing the costs of exploration.
@@
Risk-tolerant individuals thus act as functional leaders by providing information about the environment, rather than by imposing their choices on others. Their exploratory actions inform the more cautious majority, resulting in a form of passive leadership that emerges from threshold heterogeneity[King2009, Strandburg-Peshkin2018, Jolles2020], rather than social dominance or intentional coordination. This dual-speed system driven by the gradual discoveries of risk-takers coupled with rapid collective withdrawal after mass failures in foraging attempts allows for efficient collective responses without explicit coordination (Fig. 2D and SI Appendix, Fig. S3).

### v2

# Daring few, patient many: division of labor in decentralized foraging collectives

@@ Abstract
How do social animals make effective decisions in the absence of a leader? While coordination can improve accuracy, it also introduces delays as information propagates through the group. In changing environments, these delays can outweigh the benefits of globally coordinated decisions, even when local interactions remain tightly organized. This raises a key question: how can groups implement efficient collective decision-making without central coordination? We address this question using a collective foraging model in which individuals share information and rewards, but each must choose whether to bear the cost of exploring or to remain idle. We show that decentralized collectives can match the performance of centrally controlled groups through a division of labor: a small, heterogeneous subset explores even when expected rewards are negative, acquiring information to enable future foraging, while a coordinated majority forages only when expected rewards are positive. Information redundancy causes the optimal number of explorers to grow sublinearly with group size, so that larger groups need proportionally fewer explorers. The heterogeneity of the group is maximized at intermediate ecological pressures, but optimal groups are homogeneous when costs or fluctuations are extreme. Crucially, these group-level policies do not require central coordination, emerging instead from agents following simple threshold-based decision rules. We thus demonstrate a mechanism through which leaderless collectives can make effective decisions under uncertainty and show how ecological pressures can drive changes in the distribution of strategies employed by the group.
@@ 1 Introduction
One solution to this problem is centralized leadership, in which a small number of individuals guide collective action. In many wild wolf packs, for example, the breeding pair tends to initiate movements and coordinates traveling, resting, and hunting [Mech1999]. Human history offers parallel examples. Admiral Zheng He commanded fleets across the Indian Ocean in the early fifteenth century on imperial orders, spurring Chinese emigration and trade throughout Southeast Asia [Levathes2014]. A century later, royal sponsorship supported Magellan’s circumnavigation, opening the first western sea route to the Spice Islands and accelerated global trade and contact between distant regions [Bergreen2003]. In each case, a small number of designated explorers, acting under a central authority, bore enormous personal risk, while the societies that sent them reaped lasting rewards. However, most social animals, including honeybees, ants, and fish schools, lack fixed leaders and coordinate through local interactions [Pratt2002, Seeley2004, Couzin2005, Sumpter2005, Seeley2011, Hein2015]. Despite extensive empirical study of decentralized animal groups [Pratt2002, Sumpter2005, Giardina2008, Gordon2014], it remains unclear how closely decentralized coordination can match centralized control, how performance depends on diversity in individual strategies, and under what conditions heterogeneous strategies offer an advantage in leaderless groups.
@@
To address these questions, we develop and analyze a stochastic foraging model [Puterman1994, Mahadevan1996, Gold2007, Kilpatrick2021, BlumMoyse2025, Dembo2010, Bressloff2021, Barendregt2025] to understand how collectives can achieve efficient exploration. We show that well-calibrated collectives exhibit two characteristic behavioral regimes: during times of plenty, risk-averse individuals venture out together, amplifying collective gains. When resources are scarce, most individuals refrain from foraging, but a small subset of risk-tolerant explorers continues foraging to rapidly detect environmental improvements. Composition is critical as populations dominated by cautious individuals fail to discover new resources, whereas those dominated by risk-tolerant foragers exhaust shared reserves. The number of risk-tolerant individuals grows sublinearly with population size, making a balanced division of labor between exploration and exploitation essential for efficient foraging [Seeley2011, Charbonneau2015]. Remarkably, collectives with a well-calibrated mixture of cautious and bold individuals can perform as well as groups governed by a central decision-maker. The most heterogeneous compositions emerge under intermediate ecological pressures, when exploration costs and environmental variability are balanced. We thus show how simple individual decision rules lead to optimal collective performance and predict when ecological pressures favor heterogeneous versus uniform group composition.
@@
Figure 1: Overview of model and setup. Individuals in a collective observe returning foragers to update their belief about the environmental state and select actions. (A) The belief, $g^{t}$, is updated based on the observed outcomes of foraging attempts. (B) Individuals forage when their belief that the environment is in a high-reward state exceeds a threshold, $\theta$; these thresholds can differ between individuals. (C) Foraging is successful with probability $\gamma^{t}$, which switches between high and low values with the environmental state. (D) Under centralized control, a planner decides on the optimal number of foragers to maximize returns. (E) With decentralized control, coordination emerges from individual foraging decisions.
@@ 2 A Model of Collective Foraging under Uncertainty
Collective decisions about foraging in large, decentralized collectives face a fundamental challenge: how do groups balance individual risk and collective benefit in changing environments without central coordination? To address this question, we consider a minimal model designed to examine whether decentralized decision-making mechanisms can yield optimal collective long-term returns, even when individual expected returns are negative. Here we provide an overview of the model; mathematical details can be found in SI Appendix, Section 1.
@@
We assume that the community coordinates its behavior through shared information about the outcomes of foraging attempts. Returning foragers are observed by the entire community, so all individuals know how many foraging attempts have been successful (Fig. 1A), an idealization of group-level information flow in social collectives [Sumpter2005, Sumpter2008, Couzin2005]. Individuals are Bayesian observers [Wald1948, Bogacz2006, VelizCuba2016] and compute the probability that the environmental state is in a high-reward state based on the observed number of successful foragers and knowledge about the environment’s volatility. We refer to this subjective probability as their belief about the environmental state. Because all individuals have access to the same information, they share a common belief that guides their foraging decisions.
@@
Maximizing returns requires optimizing forager number: too many foragers when conditions are poor wastes effort; too few when conditions are good misses opportunities. To establish an upper bound on collective performance, we derive a theoretical benchmark based on a centralized planner that optimally determines foraging effort required to maximize long-term collective returns (Fig. 1D). Under this centralized policy, all individuals follow the directives of a planner, which infers the environmental state by observing foraging successes and then prescribes the best number of foragers. This optimal policy can be computed using dynamic programming [Bellman1954, Mahadevan1996, Sutton1998] and provides an upper bound on achievable performance. We compare this benchmark with outcomes from agents who make individual decisions based on shared information but without central coordination.
@@ 3 Results
Using our model of foraging and belief updating, we ask how a collective translates shared environmental information into effective foraging decisions. We first establish a normative benchmark based on a centralized decision-maker that maximizes expected payoff. We then show that a leaderless group can achieve optimal performance using heterogeneous individual decision thresholds and analyze how environmental conditions shape the optimal threshold distribution in a decentralized collective.
@@ 3.1 Robust collective coordination through structured heterogeneity
A central decision-maker can select an optimal number of foragers to maximize expected payoffs on each turn. In a decentralized collective, individuals share a common belief about the environmental state but make individual choices based on a private threshold. If all individuals choose a threshold that maximizes their own expected return, then all thresholds are identical. Since they share the same belief, everyone makes the same decision: forage when the expected reward is positive, and remain idle when the expected reward is negative. As a result, a collective can quickly detect when an environment switches from a high- to a low-rewarding state. However, when the collective believes the reward state is likely to be low, no individuals forage, and the collective will not detect a switch back to a high-rewarding state. The resulting failure to explore thus leads to starvation even when conditions improve. Conversely, if all individuals adopt low thresholds and forage often, the collective can quickly infer the environmental state but incurs large energetic losses in lean times. Effective collective performance therefore requires a division of labor between a minority that forages even when rewards are unlikely and a majority that waits for favorable conditions.
@@
With an appropriate distribution of individual decision thresholds, a decentralized collective can achieve the same expected returns as an optimally controlled group. We show this by constructing a threshold distribution that reproduces the number of foragers deployed under the optimal centralized policy. We do not suggest that collectives compute or implement a centralized solution; rather, this construction demonstrates that simple local threshold rules can recover centralized allocations, allowing us to analyze how ecological parameters shape the structure and robustness of the resulting strategy. The optimal strategy of a central planner can be represented by a function, $\pi(g),$ that maps the collective belief about the state of the environment, $g,$ to the number of individuals sent out to forage, $n=\pi(g)$ (Fig. 2A). As the belief that the environment is in a high-reward state increases, the planner commits more foragers [Puterman1994, Topkis1998]. The optimal policy, $\pi(g),$ is non-decreasing in the belief $g$, since higher expected rewards never justify deploying fewer foragers. We can therefore define a sequence of belief thresholds, $\theta_{n}=\inf\{\theta\in[0,1]:\pi(\theta)\geq n\}$ for $n=1,2,\cdots,N$, where $N$ is the group size. Each threshold, $\theta_{n},$ is the smallest belief value at which sending out at least $n$ foragers is optimal. Multiple thresholds may coincide, and in large groups the majority of thresholds equal the critical belief $\theta_{c}$ (Fig. 3) at which the expected foraging return is zero.
@@
Figure 2: Division of labor enables centralized and decentralized coordination. (A) In decentralized collectives, individuals have different foraging thresholds. Risk-averse individuals share a common threshold, $\theta_{c},$ and forage only when the expected reward is positive. Risk-tolerant explorers have lower, heterogeneous thresholds and forage even when the expected reward is negative. (B) Under centralized control, a planner implementing the optimal policy, $\pi(g),$ deploys an increasing number of foragers as the belief increases. An appropriate distribution of individual thresholds allows a decentralized collective to match this allocation (compare panels A and B). (C) The optimal number of risk-tolerant explorers, $N_{s},$ grows sublinearly with population size, NN, so an increasingly small fraction of explorers suffices for efficient foraging. (D) The collective belief, $g^{t},$ tracks environmental fluctuations because a subset continues exploring during lean periods, so the collective can rapidly re-engage when conditions improve.
@@
The strategy of a decentralized collective is fully determined by a set of private thresholds. Hence, a collective of size $N$ whose individuals have thresholds $\{\theta_{n}\}_{n=1}^{N}$ will send out exactly $n$ foragers when the belief satisfies $\theta_{n}\leq g<\theta_{n+1}$, with the convention $\theta_{N+1}=1$ (Fig. 2B). This choice of thresholds thus results in a number of foragers equal to that dictated by the optimal policy, $\pi(g)$ for any belief $g$. Importantly, individual decisions are based on a shared belief paired with private thresholds, rather than on orders from a central planner. The resulting distribution of foraging strategies could arise from evolved or learned differences in risk sensitivity [Seeley1983, Dreller1998, Liang2012], and is consistent with classic response-threshold models of division of labor [Bonabeau1996, Beshers2001].
@@
Collectives governed by a central decision-maker are inherently vulnerable to the loss or failure of this planner. In contrast, decentralized collectives with heterogeneous individual strategies are more robust to decision noise and environmental perturbations, potentially promoting evolutionary stability [Jandt2013, Couzin2005]. As long as phenotypic differences across the population generate an appropriate distribution of risk tolerances, the collective can achieve optimal foraging performance [Seeley1983, Charbonneau2015]. Structured heterogeneity thus allows decentralized groups to match the allocations of a central planner while maintaining robust performance in dynamic environments. We next analyze the structure of the optimal population and ask how ecological pressures shape this distribution of strategies.
3.2 Minority exploration and majority synchronization
@@
Figure 3: Effects of perturbing the distribution of decision thresholds. (A) Optimal threshold distribution for a collective of $N=100$ individuals. (B-D) Perturbed strategy distributions: (B) increasing the spread of thresholds in the risk-averse majority; (C) risk-tolerant individuals sharing a high threshold; (D) risk-tolerant individuals sharing a low threshold. Colors indicate the different perturbation types. (E) Normalized return vs. accuracy for the optimal strategy and its perturbations shows that returns are maximized at intermediate accuracy (gray dot; panel A), reflecting weakened coordinated exploitation under desynchronization (brown–gray curve; panel B), improved accuracy at increased cost under moderate heterogeneity (pink–gray curve; panel C), and degraded returns under excessive risk tolerance (blue–gray curve; panel D). Returns are normalized by the maximum possible return $\rho_{\max}$, defined as the return obtained when the entire collective forages only in the high-reward state.
@@
An optimal collective balances risk-tolerant individuals, who forage under uncertainty, with risk-averse individuals, who forage only when the expected reward is positive. Empirical studies show a consistent pattern in many social species: a small fraction of individuals, such as “scouts" in honeybee populations, persistently engages in exploration, while most colony members adopt safer strategies and exploit known resources [Seeley1983, Seeley2011, Beekman2001, Dornhaus2006]. We show that the optimal decentralized strategy yields a similar pattern. To do so, we use the critical threshold, $\theta_{c},$ and classify individuals with decision thresholds below $\theta_{c}$ as risk-tolerant, and those with thresholds equal to or exceeding $\theta_{c}$ as risk-averse.
@@
In small optimal collectives, all individuals are risk-tolerant, because each foraging outcome is highly informative and has a large impact on the collective belief about the environmental state (Fig. 2C). As group size increases, performance improves, but the optimal division of labor changes. The expected cost of foraging grows linearly with the number of active foragers, whereas additional foraging outcomes provide redundant information and thus become less valuable. A mathematical analysis shows that the optimal number of risk-tolerant individuals grows logarithmically with group size (dashed line in Fig. 2C), reflecting diminishing informational returns from additional explorers as observations become increasingly redundant. This sublinear scaling is robust across a wide range of environmental volatilities, foraging costs, and reward structures (SI Appendix, Fig. S1). Consequently, larger groups comprise a growing majority of risk-averse individuals, while the fraction of risk-tolerant explorers declines. This prediction is consistent with empirical and theoretical studies showing that effective animal collectives rely on a small exploratory minority whose size grows more slowly than group size [Couzin2005, Cronin2014]. Details of the asymptotic analysis underlying this result are provided in SI Appendix, Section 2.
@@
While the decision thresholds of risk-tolerant individuals are distributed, all risk-averse individuals share a common threshold, $\theta_{c}$. Indeed, the entire collective commits to foraging once the expected reward becomes positive (see SI Appendix and Fig. S2). This synchronization of foraging bouts is a consequence of the linear increase of total reward with the number of committed individuals, implying that full commitment is optimal when returns are expected to be favorable. Thus, an optimal collective consists of a heterogeneous minority that maintains environmental awareness through continued sampling, and a synchronized majority that maximizes collective returns through coordinated exploitation once conditions are deemed favorable.
@@
Optimal collective foraging strategies exhibit a characteristic cascade of activity in response to environmental changes (Fig. 2D). When conditions deteriorate, widespread foraging produces few successes, quickly shifting the collective belief toward low expected rewards. Most individuals stop foraging, leaving a small subset of risk-tolerant explorers that continue sampling. When conditions deteriorate severely, even limited exploration ceases to minimize costs. Nevertheless, in the absence of new information, collective belief gradually relaxes toward uncertainty, prompting renewed exploration. If some of these sparse foraging attempts are successful, the belief shifts towards high expected rewards, and more foragers are recruited. Once the collective belief crosses the threshold $\theta_{c}$, coordinated collective foraging is triggered again, and the process repeats. Heterogeneity in decision thresholds enables the population to function as a distributed sensor, maintaining sensitivity to environmental changes while balancing the cost of exploration.
@@
Risk-tolerant individuals thus act as functional leaders by providing information about the environment, rather than by imposing their choices on others. Their exploratory actions inform the more cautious majority, resulting in a form of passive leadership that emerges from threshold heterogeneity[King2009, Strandburg-Peshkin2018, Jolles2020], rather than social dominance or intentional coordination. This dual-speed system, driven by the gradual discoveries of risk-takers coupled with rapid collective withdrawal after mass failures in foraging attempts, allows for efficient collective responses without explicit coordination (Fig. 2D and SI Appendix, Fig. S3).